# Variational Inference for Sparse Poisson Regression

### v8

@@ 5 Posterior predictive mass function
The predictive probability mass function in (18) can be obtained using a univariate numerical integration, and might be used for obtaining a point predictor (e.g., most likely predictor) and a prediction interval (e.g., the most likely prediction interval) for $y_{0}$ given $x_{0}$ and $(X,y)$.
@@ 6 Simulation study
To examine the performance of the proposed models, we have conducted a simulation study as follows. In each of $N=1000$ replications, we have generated samples of size $n$, from a sparse Poisson regression model, whose vector of coefficients ${\boldsymbol{\beta}}_{p\times 1}$, with $p$, is first generated from $N(\mu_{0},\sigma_{0})$ and then is multiplied by a vector $\mathbf{z}$ of length $p$ with 0 and 1 elements, to set some elements of $\boldsymbol{\beta}$} to zero. The rows of the covariate matrix $X$ are generated independently from a multivariate normal distribution with mean $\mu_{X}$ and a variance-covariance matrix with elements ${\Sigma}_{ij}=\sigma^{2}_{X}\cdot 0.3^{|i-j|}$, to model a slight multicollinearity. We consider the cases $n=30,100$, and $p=10,200$, to successfully cover the low- and high-dimensional scenarios. For the cases with $p=10$, we let $\mu_{0}=0.7,\sigma_{0}=0.5,\mu_{X}=0.1,\sigma^{2}_{X}=1$, and we set $\mathbf{z}=(1,0,1,0,0,0,1,0,1,0)^{\top}$, and for the cases with $p=200$, we let $\mu_{0}=0.1,\sigma_{0}=0.6,\mu_{X}=0.1,\sigma^{2}_{X}=0.05$, and we randomly choose 60 (including the first one) out of 200 values of $\mathbf{z}$ to be equal 1, and the remaining are set to 0. The case n=100,p=10n=100,\;p=10, called the low-dimensional scenario, and the case n=30,p=200n=30,\;p=200, called the high-dimensional scenario are studied in the following subsections, while the two remaining cases (n=30,p=10n=30,\;p=10 and n=100,p=200n=100,\;p=200) are given in the supplementary material for the sake of brevity.
@@
where ${y}_{t}^{\rm train}=({y}_{1t},\ldots,{y}_{nt})^{\top}$ and ${y}_{t}^{\rm test}=({y}^{\prime}_{1t},\ldots,{y}^{\prime}_{n^{\prime}t})^{\top}$, are the train-set (80% of the full sample) and test-set (20 % of the full sample) response samples, respectively, generated during $t$th iteration.
@@ 6.1 Low dimensional scenario
Figures 6 and 7 show the box plots of all aforementioned criteria for all competitive methods, for the low-dimensional scenario, with $p=10$, $n=100$, and $\mathbf{z}=(1,0,1,0,1)^{\top}$. In this scenario, all models are also compared to their MCMC versions. In implementing the MCMC methods, we utilized the R packages R2OpenBUGS and rjags within R (version 4.2.1), in conjunction with OpenBUGS (version 3.2.3) and JAGS (version 4.3.1). Each chain comprised 10,000 iterations, with a burn-in of 5,000 and a thinning interval of 10.
Among the VB variants, Bernoulli–VB achieve the lowest median CRE with limited variability, whereas CS-VB and LAPLACE-VB show slightly more dispersion and occasional outliers, suggesting sensitivity to data characteristics or prior assumptions. Comparisons with the corresponding MCMC counterparts (Bernoulli-MCMC, CS-MCMC, and LAPLACE-MCMC) indicate that the VB framework can achieve similar accuracy levels at a markedly reduced computational cost, albeit with modest trade-offs in variance. Overall, the figure underscores the proposed VB methods as effective and computationally efficient alternatives for coefficient estimation, maintaining accuracy close to established methods while offering scalability advantages.
@@
All methods achieve relatively low median TSRE values (around 0.05-0.08), suggesting strong predictive performance. The VB methods are competitive with established baselines, with Bernoulli-VB and LAPLACE-VB displaying compact interquartile ranges comparable to LASSO and SCAD, while CS-VB shows a slightly higher median and a few large-value outliers, potentially reflective of sensitivity to specific test scenarios. Compared to their MCMC counterparts, the VB methods maintain similar median accuracy but generally exhibit fewer extremely high errors, underscoring a balance between stability and efficiency. Notably, the sparsity-inducing VB variants (CS-VB and LAPLACE-VB) deliver consistent out-of-sample performance while avoiding the heavier computational cost of MCMC sampling, reinforcing their practicality for large-scale predictive modeling. Overall, the figure highlights the ability of the proposed VB framework to retain robust generalization while ensuring computational scalability.
@@
Although the computation time of the VB methods is a little more than the LASSO and SCAD methods, this comparison is not fair, since the LASSO and SCAD methods use C++ programming for accelerating the computation, while we have done all computations in the R software. So, it seems that the VB methods are very fast, as expected. The CS-VB method seems to be the fastest VB method.
Table 1 presents the average (standard deviation) of the coverage probabilities of the highest posterior density (HPD) confidence intervals for the regression coefficients for three VB methods at level 0.95. Table 1 indicates that the average coverage probabilities of nearly all regression coefficients approximate the nominal value.
@@
Figure 5: The boxplots of the accuracies for the resgression coefficients for different VB-methods.
@@
Figure 7: The low-dimensional scenario simulation results: the FPR (top), the TPR (middle), and the relative computation time (bottom) for different methods.
@@
${\rm accuracy}(q)=100\left(1-\frac{1}{2}\int_{-\infty}^{\infty}|q(\theta_{j})-p(\theta_{j}|y)|d\theta\right)\%,$ (21)
@@
where $p(\theta_{j}|y)$ is the kernel density estimator using the MCMC sample. If $q(\theta)=p(\theta|y)$ for all values of $\theta$, we have ${\rm accuracy}(q)=100\%,$ while since both $q(\cdot)$ and $p(\cdot|y)$ are probability density functions, the minimum value of ${\rm accuracy}(q)$ is zero, which occurs when $q(\cdot)$ and $p(\cdot|y)$ have separate supports. For the Bernoulli components $q(\gamma_{j})$ and $q(Z_{j})$ for Bernoulli-VB and CS-VB methods, the accuracy is computed as follows
${\rm accuracy}(q)=100\left(1-\frac{1}{2}\left(\left|q(\theta_{j}=1)-\frac{1}{T}\sum_{t=1}^{T}I(\theta_{jt}=1)\right|+\left|q(\theta_{j}=0)-\frac{1}{T}\sum_{t=1}^{T}I(\theta_{jt}=0)\right|\right)\right)\%,$
@@
where $\theta_{j1},\ldots,\theta_{jT}$ are MCMC samples. If $q({\rho}_{j}=k)=\frac{1}{T}\sum_{t=1}^{T}I({\rho}_{jt}=k)$ for $k=0,1$, we have ${\rm accuracy}(q)=100\%,$ and when $q({\rho}_{j}=k)=1$ and $\frac{1}{T}\sum_{t=1}^{T}I({\rho}_{jt}=k)=0$ for some $k\in\{0,1\}$, we have ${\rm accuracy}(q)=0\%$.
@@
The average (standard error) of accuracy values are given in Table 2. The boxplots for the accuracies of the regression coefficients are given in Figure 5. Furthermore, the $q$-functions as well as the kernel density estimator of the marginal posteriors of the regression coefficients, for a single iteration of the simulation study, are given in Figures 8, 9, and the corresponding plots for the other parameters are given in the supplementary material.
@@
As one can see from Table 2 and Figure 5, the accuracy of the posterior approximation for the regression coefficients is high, except for zero coefficients in the Bernoulli-VB method. Note that the variational posteriors of the regression coefficients under the Bernoulli-VB method are not directly comparable to their MCMC estimated densities. In this case, we would need to compute the accuracy of the variational posterior of ${\boldsymbol{\beta}}\Gamma$. Thus, the marginal densities are either $q(\beta_{j})$ or degenerate at zero (when $\gamma_{j}=0$). Based on the criterion in Equation (21), the accuracy of the variational posterior is reduced, while in this scenario the point mass at zero offers a notably more accurate posterior. Table 2 also presents the accuracy for the parameters $\boldsymbol{\gamma}$} and $\mathbf{Z}$ of the VB methods against an MCMC benchmark for the Bernoulli and CS priors. As can be seen from these values, the CS-VB has higher accuracies than the Bernoulli-VB method. The accuracy for the $\pi$ hyperparameter is moderate for both Bernoulli-VB and CS-VB methods. Regarding the variance hyper-parameters of the Bernoulli-VB and Laplace-VB methods, we can see from Table 2 that the Bernoulli-VB method has lower accuracies for the variance components, which correspond with the zero coefficients compared to the other components, while the Laplace-VB method achieves almost equal accuracies for all components $\tau_{0},\ldots,\tau_{4}$.
@@
Figure 12: The high-dimensional scenario simulation results: the FPR (top), the TPR (middle), and the computation time (bottom) for different methods.
6.2 High dimesional scenario
To evaluate the performance of the proposed methods in a high-dimensional setting, we consider the case with $p=200$, $n=30$, and We randomly selected 60 values of $\mathbf{z}$ (including the first) to be set to 1, with the remaining values assigned 0.. Figures 11 and 12 show the boxplots of all aforementioned criteria for all competitive methods.
@@ 7 Benchmark real data analysis
To examine the prediction performance of the proposed VB methods, we have considered 5 benchmark real data sets with a count response variable. The affairs data set (Fair, 1978), available in the R package COUNT, contains 601 observations and 18 variables, including naffairs (the number of affairs in the past year) and 17 covariates related to children, marital happiness, religiosity, and age. Greene (2003), modeled this data using Poisson regression, although given the amount of over-dispersion in the data, employing a negative binomial model is an appropriate strategy. The variable naffairs is considered the response variable. The top left graph of Figure 13 shows the histogram of the response variable naffairs. The bike-sharing dataset records hourly and daily rental counts from 2011 to 2012 in the Capital Bikeshare system, along with corresponding weather and seasonal information. This data set includes 731 observations and 14 variables (except date and instant number), including cnt (the response variable), and 13 covariates, including season, year, month, holiday, weekday, working day, 5 weather variables, casual, and registered. The top middle graph in Figure 13 presents the histogram of the response variable for the bike-sharing data set. The three remaining data sets, azcabgptca, azdrg112, and azpro, from the R package COUNT pertain to samples from Arizona hospital cardiovascular patient files, collected in 1991 (azcabgptca, azpro) and 1995 (azdrg112), involving patients who received one of two standard procedures: CABG or PTCA. The data set azcabgptca has 1959 observations on 6 variables, azdrg112 has 1,798 observations on 4 variables, and azpro has 3589 observations on 6 variables. In all three data sets, the response variable is los (length of hospital stay), with covariates including procedure type, sex, age, and additional factors. The bottom plots of Figure 13 show the histograms of the los response variable in azcabgptca, azdrg112, and azpro, respectively, from left to right.

### v9

@@ 4 Sparsity-enforcing thresholds
The AIC criterion is asymptotically equivalent to leave-one-out cross-validation and is specifically designed to find the model with the best predictive performance. We have preferred AIC rather than BIC, since the BIC criterion is less effective for prediction in finite samples, especially when the true model is complex and may not be in the candidate set. BIC tends to select more parsimonious models than AIC, which can lead to underfitting and poorer predictive performance. For the hard thresholding procedure, when the dimension is high, we need to evaluate a large number of potential models. Using AIC provided a computationally efficient and deterministic criterion without the need for extensive resampling, which would have been computationally prohibitive.
The above hard-thresholding strategy is a suitable choice for our framework compared to soft-thresholding and posterior inclusion probabilities. The soft thresholding produces sparsity via continuous shrinkage, which biases the non-zero coefficients toward zero. Furthermore, the posterior inclusion probabilities from the VB approximation are poorly calibrated due to the known over-confidence of the mean-field approximation, making them unreliable as a soft selection tool. Therefore, applying a hard threshold to the coefficients is the most coherent way to distill a single, interpretable model from the approximated posterior, providing a clear set of selected variables without additional shrinkage.
@@ 5 Posterior predictive mass function
The predictive probability mass function in (18) can be obtained using a univariate numerical integration, and might be used for obtaining a point predictor (e.g., most likely predictor) and a prediction interval (e.g., the most likely prediction interval) for $y_{0}$ given $x_{0}$ and $(X,y)$. Evaluating this single-variable integral with deterministic numerical methods (e.g., Gaussian quadrature) is computationally very cheap and achieves high precision without the sampling error inherent in Monte Carlo methods, which would require drawing a large number of samples from a high-dimensional multivariate posterior. Consequently, the above method is not only more accurate but also faster than a sampling-based alternative for calculating prediction intervals.
@@ 6 Simulation study
To examine the performance of the proposed models, we have conducted a simulation study as follows. In each of $N=1000$ replications, we have generated samples of size $n$, from a sparse Poisson regression model, whose vector of coefficients ${\boldsymbol{\beta}}_{p\times 1}$, with $p$, is first generated from $N(\mu_{0},\sigma_{0})$ and then is multiplied by a vector $\mathbf{z}$ of length $p$ with 0 and 1 elements, to set some elements of $\boldsymbol{\beta}$} to zero. The rows of the covariate matrix $X$ are generated independently from a multivariate normal distribution with mean $\mu_{X}$ and a variance-covariance matrix with elements ${\Sigma}_{ij}=\sigma^{2}_{X}\cdot 0.3^{|i-j|}$, to model a slight multicollinearity. We consider the cases $n=30,100$, and $p=10,200$, to successfully cover the low- and high-dimensional scenarios. For the cases with $p=10$, we let $\mu_{0}=0.7,\sigma_{0}=0.5,\mu_{X}=0.1,\sigma^{2}_{X}=1$, and we set $\mathbf{z}=(1,0,1,0,0,0,1,0,1,0)^{\top}$, and for the cases with $p=200$, we let $\mu_{0}=0.1,\sigma_{0}=0.6,\mu_{X}=0.1,\sigma^{2}_{X}=0.05$, and we randomly choose 60 (including the first one) out of 200 values of $\mathbf{z}$ to be equal 1, and the remaining are set to 0. The case n=100,p=10n=100,\;p=10, called the low-dimensional scenario, and the case n=30,p=200n=30,\;p=200, called the high-dimensional scenario, are studied in the following subsections, while the two remaining cases (n=30,p=10n=30,\;p=10 and n=100,p=200n=100,\;p=200) are given in the supplementary material for the sake of brevity.
@@
where ${y}_{t}^{\rm train}=({y}_{1t},\ldots,{y}_{nt})^{\top}$ and ${y}_{t}^{\rm test}=({y}^{\prime}_{1t},\ldots,{y}^{\prime}_{n^{\prime}t})^{\top}$, are the train-set (80% of the full sample) and test-set (20 % of the full sample) response samples, respectively, generated during $t$th iteration, $\bar{y}_{t}^{\rm train}=\frac{1}{n}\sum_{i=1}^{n}{y}_{it}$ and $\bar{y}_{t}^{\rm test}=\frac{1}{n^{\prime}}\sum_{i=1}^{n}{y^{\prime}}_{it}$.
@@ 6.1 Low dimensional scenario
Figures 7 and 8 show the box plots of all aforementioned criteria for all competitive methods, for the low-dimensional scenario, with $p=10$, $n=100$, and $\mathbf{z}=(1,0,1,0,1)^{\top}$. In this scenario, all models are also compared to their MCMC versions. In implementing the MCMC methods, we utilized the R packages R2OpenBUGS and rjags within R (version 4.2.1), in conjunction with OpenBUGS (version 3.2.3) and JAGS (version 4.3.1). Each chain comprised 10,000 iterations, with a burn-in of 5,000 and a thinning interval of 10.
Among the VB variants, Bernoulli-VB achieve the lowest median CRE with limited variability, whereas CS-VB and LAPLACE-VB show slightly more dispersion and occasional outliers, suggesting sensitivity to data characteristics or prior assumptions. Comparisons with the corresponding MCMC counterparts (Bernoulli-MCMC, CS-MCMC, and LAPLACE-MCMC) indicate that the VB framework can achieve similar accuracy levels at a markedly reduced computational cost, albeit with modest trade-offs in variance. Overall, the figure underscores the proposed VB methods as effective and computationally efficient alternatives for coefficient estimation, maintaining accuracy close to established methods while offering scalability advantages.
@@
All methods achieve relatively low median TSRE values (around 0.05-0.08), suggesting strong predictive performance. The VB methods are competitive with established baselines, with Bernoulli-VB and LAPLACE-VB displaying compact interquartile ranges comparable to LASSO and SCAD, while CS-VB shows a slightly higher median and a few large-value outliers, potentially reflective of sensitivity to specific test scenarios. Compared to their MCMC counterparts, the VB methods maintain similar median accuracy but generally exhibit fewer extremely high errors, underscoring a balance between stability and efficiency. Notably, the sparsity-inducing VB variants (CS-VB and LAPLACE-VB) deliver consistent out-of-sample performance while avoiding the higher computational cost of MCMC sampling, reinforcing their practicality for large-scale predictive modeling. Overall, the figure highlights the ability of the proposed VB framework to retain robust generalization while ensuring computational scalability.
@@
Although the computation time of the VB methods is a little more than the LASSO and SCAD methods, this comparison is not fair, since the LASSO and SCAD methods use C++ programming for accelerating the computation, while we have done all computations in the R software. So, it seems that the VB methods are very fast, as expected. The CS-VB method seems to be the fastest VB method. All computations were performed using R version 4.3.2 on a machine with a Core i5-10210U CPU.
Table 1 presents the average (standard deviation) of the coverage probabilities of the highest posterior density (HPD) confidence intervals for the regression coefficients for three VB methods at level 0.95. Table 1 indicates that the average coverage probabilities of nearly all regression coefficients approximate the nominal value. In some cases, the average coverage probabilities for the HPD intervals are occasionally below the nominal 0.95 level for both the VB and MCMC methods. This pattern is observed even for the MCMC benchmarks, which suggests that the under-coverage is not a consequence of the variational approximation but is inherent to the model and prior specification. It reflects the frequentist operating characteristics of the Bayesian procedure with these specific priors in the given simulation setting. The key finding is that the VB methods closely replicate the coverage properties of the corresponding MCMC methods, demonstrating their reliability for uncertainty quantification in this context.
@@
Figure 6: The boxplots of the accuracies for the regression coefficients for different VB-methods.
@@
Figure 8: The low-dimensional scenario simulation results: the FPR (top), the FNR (middle), and the relative computation time (bottom) for different methods.
@@
${\rm accuracy}(q(\theta_{j}))=100\left(1-\frac{1}{2}\int_{-\infty}^{\infty}|q(\theta_{j})-p(\theta_{j}|y)|d\theta\right)\%,$ (21)
@@
where $p(\theta_{j}|y)$ is the kernel density estimator using the MCMC sample. If $q(\theta)=p(\theta|y)$ for all values of $\theta$, we have ${\rm accuracy}(q(\theta_{j}))=100\%,$ while since both $q(\cdot)$ and $p(\cdot|y)$ are probability density functions, the minimum value of ${\rm accuracy}(q(\theta_{j}))$ is zero, which occurs when $q(\cdot)$ and $p(\cdot|y)$ have separate supports. For the Bernoulli components $q(\gamma_{j})$ and $q(Z_{j})$ for Bernoulli-VB and CS-VB methods, the accuracy is computed as follows
${\rm accuracy}(q(\theta_{j}))=100\left(1-\frac{1}{2}\left(\left|q(\theta_{j}=1)-\frac{1}{T}\sum_{t=1}^{T}I(\theta_{jt}=1)\right|+\left|q(\theta_{j}=0)-\frac{1}{T}\sum_{t=1}^{T}I(\theta_{jt}=0)\right|\right)\right)\%,$
@@
where $\theta_{j1},\ldots,\theta_{jT}$ are MCMC samples. If $q({\theta}_{j}=k)=\frac{1}{T}\sum_{t=1}^{T}I({\theta}_{jt}=k)$ for $k=0,1$, we have ${\rm accuracy}(q(\theta_{j}))=100\%,$ and when $q({\theta}_{j}=k)=1$ and $\frac{1}{T}\sum_{t=1}^{T}I({\theta}_{jt}=k)=0$ for some $k\in\{0,1\}$, we have ${\rm accuracy}(q(\theta_{j}))=0\%$.
@@
The average (standard error) of accuracy values are given in Table 2. The boxplots for the accuracy of the regression coefficients are given in Figure 6. Furthermore, the $q$-functions as well as the kernel density estimator of the marginal posteriors of the regression coefficients, for a single iteration of the simulation study, are given in Figures 9, 10, and the corresponding plots for the other parameters are given in the supplementary material.
@@
As one can see from Table 2 and Figure 6, the accuracy of the posterior approximation for the regression coefficients is high, except for zero coefficients in the Bernoulli-VB method. Note that the variational posteriors of the regression coefficients under the Bernoulli-VB method are not directly comparable to their MCMC estimated densities. In this case, we would need to compute the accuracy of the variational posterior of ${\boldsymbol{\beta}}\Gamma$. Thus, the marginal densities are either $q(\beta_{j})$ or degenerate at zero (when $\gamma_{j}=0$). Based on the criterion in Equation (21), the accuracy of the variational posterior is reduced, while in this scenario the point mass at zero offers a notably more accurate posterior. For the simulation study where the true $\beta_{j}$ is known, if we replace $p(\beta_{j}|y)$ by $\delta(\beta_{j})p(\beta_{j}|y)$ in (21), for the Bernoulli-VB method, we obtain an average accuracy of 82.89, 75.00, 82.89, 82.89, 94.74, and 82.89, for $\beta_{1},\beta_{3},\beta_{4},\beta_{5},\beta_{7}$, and $\beta_{9}$, respectively.
Table 2 also presents the accuracy for the parameters $\boldsymbol{\gamma}$} and $\mathbf{Z}$ of the VB methods against an MCMC benchmark for the Bernoulli and CS priors. As can be seen from these values, the CS-VB has higher accuracies than the Bernoulli-VB method. The accuracy for the $\pi$ hyperparameter is moderate for both Bernoulli-VB and CS-VB methods. Regarding the variance hyper-parameters of the Bernoulli-VB and Laplace-VB methods, we can see from Table 2 that the Bernoulli-VB method has lower accuracies for the variance components, which correspond with the zero coefficients compared to the other components, while the Laplace-VB method achieves almost equal accuracies for all components $\tau_{0},\ldots,\tau_{4}$.
@@
Figure 13: The high-dimensional scenario simulation results: the FPR (top), the FNR (middle), and the computation time (bottom) for different methods.
6.2 High dimensional scenario
To evaluate the performance of the proposed methods in a high-dimensional setting, we consider the case with $p=200$, $n=30$, and We randomly selected 60 values of $\mathbf{z}$ (including the first) to be set to 1, with the remaining values assigned 0.. Figures 12 and 13 show the boxplots of all aforementioned criteria for all competitive methods.
@@ 7 Benchmark real data analysis
To examine the prediction performance of the proposed VB methods, we have considered 5 benchmark real data sets with a count response variable. It is worth noting that we do not claim that these datasets are suitable for Poisson regression. Indeed, all these datasets exhibit overdispersion, making a Negative Binomial model a strong candidate; however, our primary aim is to benchmark the performance of the proposed sparse variational Bayes algorithms within the Poisson framework. Using these challenging, real-world datasets provides a stringent test that demonstrates our method’s robustness in accurate variable selection and its computational advantages, even under realistic conditions of model misspecification.
The affairs data set (Fair, 1978), available in the R package COUNT, contains 601 observations and 18 variables, including naffairs (the number of affairs in the past year) and 17 covariates related to children, marital happiness, religiosity, and age. Greene (2003), modeled this data using Poisson regression, although given the amount of over-dispersion in the data, employing a negative binomial model is an appropriate strategy. The variable naffairs is considered the response variable. The top left graph of Figure 14 shows the histogram of the response variable naffairs. The bike-sharing dataset records hourly and daily rental counts from 2011 to 2012 in the Capital Bikeshare system, along with corresponding weather and seasonal information. This data set includes 731 observations and 14 variables (except date and instant number), including cnt (the response variable), and 13 covariates, including season, year, month, holiday, weekday, working day, 5 weather variables, casual, and registered. The top middle graph in Figure 14 presents the histogram of the response variable for the bike-sharing data set. The three remaining data sets, azcabgptca, azdrg112, and azpro, from the R package COUNT pertain to samples from Arizona hospital cardiovascular patient files, collected in 1991 (azcabgptca, azpro) and 1995 (azdrg112), involving patients who received one of two standard procedures: CABG or PTCA. The data set azcabgptca has 1959 observations on 6 variables, azdrg112 has 1,798 observations on 4 variables, and azpro has 3589 observations on 6 variables. In all three data sets, the response variable is los (length of hospital stay), with covariates including procedure type, sex, age, and additional factors. The bottom plots of Figure 14 show the histograms of the los response variable in azcabgptca, azdrg112, and azpro, respectively, from left to right.